# Two-Step Regularized HARX to Measure Volatility Spillovers in Multi-Dimensional Systems

### v2

@@ 4.2 Spillover Network Structure
Figure 3: Hybrid HAR-ElasticNet Coefficient Matrix. Notes: Rows show the 18 features (source asset $\times$ frequency: daily, weekly average, monthly average). Columns show the 6 target assets. Each cell displays the estimated coefficient. Diagonal blocks (e.g., ZS features predicting ZS) show own-market persistence from the OLS HAR step, preserving realistic  0.99 persistence. Off-diagonal cells show cross-market spillovers from the ElasticNet step. White or near-zero cells indicate coefficients shrunk to zero by regularization. The hybrid approach yields extreme sparsity in cross-market terms: only 7 of 90 cross-market coefficients (8%) are nonzero, with ES, NQ, ZF, and ZN having zero cross-market spillovers, while ZS and CL retain sparse transmission channels.
@@ 4.3 Impulse Response Analysis: Commodities Shock
The commodity shock generates strong, persistent within-group effects. Crude oil volatility increases by approximately 0.17 (roughly 50% relative to its unconditional mean of 0.34) on impact, decaying slowly to 0.12 by day 20—retaining approximately 71% of the initial shock. This slow decay reflects the high own-lag persistence (around 0.99) preserved by the hybrid model’s OLS HAR step. Soybean volatility shows a smaller initial response of approximately 0.08 (about 13% of its unconditional mean of 0.61), decaying to 0.05 by day 20 (62% of impact remains).
@@ 4.5 Impulse Response Analysis: Treasuries Shock
Treasury shocks exhibit realistic persistence within the treasury complex but no cross-market transmission making it to the mean response. Both 5-year and 10-year treasury volatility show impact responses near zero, as expected given their low unconditional volatility (mean RV of 0.04 for both). While the absolute magnitudes are small, the relative persistence is consistent with other assets: treasury volatility decays slowly, retaining the characteristic roughly 0.99 persistence from the HAR own-lag structure. Cross-market spillovers are completely absent, but soybeans and crude oil show an increase in volatility of volatility in that the width of the 95% confidence intervals grow over the reponse horizon.
@@ 5 Conclusion
This paper demonstrates that separating own-volatility dynamics from cross-market spillovers is essential for credible network identification. Using over 20 years of daily realized volatility across six futures markets, we develop a hybrid HAR-ElasticNet framework that estimates own-lag persistence via OLS (preserving the characteristic roughly 0.99 persistence of volatility) before applying ElasticNet regularization to cross-market terms. This two-step procedure reveals an extremely sparse network: only 7 of 90 cross-market spillover coefficients (8%) are nonzero, with equities (ES, NQ) and treasuries (ZF, ZN) exhibiting zero statistically significant cross-market spillovers. The surviving spillovers are concentrated in commodity markets, where crude oil receives small spillovers from equities and treasuries, while soybeans remain largely isolated. We did observe some response in the volatility of volatility as crude oil and soybean JIRF confidence intervals grew over the response horizons when equity or treasuries were shocked.
This extreme sparsity contrasts sharply with standard HARX-ElasticNet, which identifies 40-50% nonzero cross-market coefficients and suggests equities as dominant transmitters. Our methodological contribution is to show that this discrepancy reflects a fundamental issue that must be delt with to properly identify spillover networks. Uniform regularization applied to all coefficients (including own-lags) shrinks persistence from 0.99 to 0.08-0.18, forcing the model to compensate by attributing persistence to spurious cross-market pathways. The hybrid approach corrects this mis-specification, revealing that volatility is predominantly a within-market phenomenon.

### v3

@@ 4.2 Spillover Network Structure
Figure 3: Hybrid HARX-ElasticNet Coefficient Matrix. Notes: Rows show the 18 features (source asset $\times$ frequency: daily, weekly average, monthly average). Columns show the 6 target assets. Each cell displays the estimated coefficient. Diagonal blocks (e.g., ZS features predicting ZS) show own-market persistence from the OLS HAR step, preserving realistic  0.99 persistence. Off-diagonal cells show cross-market spillovers from the ElasticNet step. White or near-zero cells indicate coefficients shrunk to zero by regularization. The hybrid approach yields extreme sparsity in cross-market terms: only 7 of 90 cross-market coefficients (8%) are nonzero, with ES, NQ, ZF, and ZN having zero cross-market spillovers, while ZS and CL retain sparse transmission channels.
@@ 4.3 Impulse Response Analysis: Commodities Shock
The commodity shock generates strong, persistent within-group effects. Crude oil volatility increases by approximately 0.17 (roughly 50% relative to its unconditional mean of 0.34) on impact, decaying slowly to 0.12 by day 20—retaining approximately 71% of the initial shock. This slow decay reflects the high own-lag persistence (around 0.99) preserved by the hybrid model’s OLS HAR step. Soybean volatility shows a smaller initial response of approximately 0.08 (about 39% of its unconditional mean of 0.207), decaying to 0.05 by day 20 (62% of impact remains).
@@ 4.5 Impulse Response Analysis: Treasuries Shock
Treasury shocks exhibit realistic persistence within the treasury complex but no cross-market transmission making it to the mean response. Both 5-year and 10-year treasury volatility show impact responses near zero, as expected given their low unconditional volatility (mean RV of 0.04 for both). While the absolute magnitudes are small, the relative persistence is consistent with other assets: treasury volatility decays slowly, retaining the characteristic roughly 0.99 persistence from the HAR own-lag structure. Cross-market spillovers are completely absent, but soybeans and crude oil show an increase in volatility of volatility in that the width of the 95% confidence intervals grow over the response horizon.
@@ 5 Conclusion
This paper demonstrates that separating own-volatility dynamics from cross-market spillovers is essential for credible network identification. Using over 20 years of daily realized volatility across six futures markets, we develop a hybrid HARX-ElasticNet framework that estimates own-lag persistence via OLS (preserving the characteristic roughly 0.99 persistence of volatility) before applying ElasticNet regularization to cross-market terms. This two-step procedure reveals an extremely sparse network: only 7 of 90 cross-market spillover coefficients (8%) are nonzero, with equities (ES, NQ) and treasuries (ZF, ZN) exhibiting zero statistically significant cross-market spillovers. The surviving spillovers are concentrated in commodity markets, where crude oil receives small spillovers from equities and treasuries, while soybeans remain largely isolated. We did observe some response in the volatility of volatility as crude oil and soybean JIRF confidence intervals grew over the response horizons when equity or treasuries were shocked.
This extreme sparsity contrasts sharply with standard HARX-ElasticNet, which identifies 40-50% nonzero cross-market coefficients and suggests equities as dominant transmitters. Our methodological contribution is to show that this discrepancy reflects a fundamental issue that must be dealt with to properly identify spillover networks. Uniform regularization applied to all coefficients (including own-lags) shrinks persistence from 0.99 to 0.08-0.18, forcing the model to compensate by attributing persistence to spurious cross-market pathways. The hybrid approach corrects this mis-specification, revealing that volatility is predominantly a within-market phenomenon.